OpenStudy (anonymous):

What is the amplitude of y = 1/2 sin 2x?

OpenStudy (kainui):

What is the amplitude of \(\sin(2x)\)? Do you see why the amplitude of \(\frac{1}{2}\sin(2x)\) would exactly be the half of the amplitude of \(\sin(2x)\)?

OpenStudy (anonymous):

im lost you confused me

OpenStudy (anonymous):

1?